Three years ago, Debra King started a business that creates and delivers holiday and birthday gift baskets to students at the local university. Debra sells the baskets for $34 each, and her variable costs are $24 per basket. She incurs $24,000 in fixed costs each year.

Last year, Debra sold 5,000 baskets, and she believes that demand this year will be stable at 5,000 baskets. The following are the actions Debra could take if she wants to earn $31,000 in operating income by selling only 5,000 baskets. Consider each action independently. (Round per unit answers to 2 decimal places, e.g. 52.75 and fixed cost to 0 decimal places, e.g. 5,275.)

1.

Raise selling price per unit to

$enter raise in selling price per unit amount in dollar rounded to 2 decimal places

2.

Reduce variable costs per unit to

$enter reduce variable costs per unit amount in dollar rounded to 2 decimal places

3.

Reduce fixed costs to

$enter reduce fixed costs per unit amount in dollar rounded to 2 decimal places
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Debra King
Answer 1 Amount $ Note
Sell price           34.00 A
Less: Variable cost           24.00 B
Contribution per basket           10.00 C
Number of baskets      5,000.00 D
Total Contribution 50,000.00 E=C*D
Less: Fixed costs 24,000.00 F
Current Operating income 26,000.00 G
Desired Operating income 31,000.00 H
Difference     5,000.00 I=H-G
Difference per unit             1.00 K=I/D
Revised Sell price           35.00 L=K+A
Answer 2 Amount $ Note
Sell price           34.00 A
Less: Variable cost           24.00 B
Contribution per basket           10.00 C
Number of baskets      5,000.00 D
Total Contribution 50,000.00 E=C*D
Less: Fixed costs 24,000.00 F
Current Operating income 26,000.00 G
Desired Operating income 31,000.00 H
Difference     5,000.00 I=H-G
Difference per unit             1.00 K=I/D
Revised Variable cost           23.00 L=B-K
Answer 3 Amount $ Note
Sell price           34.00 A
Less: Variable cost           24.00 B
Contribution per basket           10.00 C
Number of baskets      5,000.00 D
Total Contribution 50,000.00 E=C*D
Less: Fixed costs 24,000.00 F
Current Operating income 26,000.00 G
Desired Operating income 31,000.00 H
Difference     5,000.00 I=H-G
Revised Fixed cost 19,000.00 J=F-I